# LLM Agent Honeypot: Monitoring AI Hacking Agents in the Wild https://ai-honeypot.palisaderesearch.org

### v1

@@ Abstract
We introduce the LLM Honeypot, a system for monitoring autonomous AI hacking agents. We deployed a customized SSH honeypot and applied prompt injections with temporal analysis to identify LLM-based agents among attackers. Over a trial run of a few weeks in a public environment, we collected 800,000 hacking attempts and 6 potential AI agents, which we plan to analyze in depth in future work. Our objectives aim to improve awareness of AI hacking agents and enhance preparedness for their risks.
@@ 1 Introduction
The continuous evolution of AI capabilities and agent frameworks is gradually increasing the potential for AI-driven cyberattacks. These advancements make it possible to create autonomous agents capable of adapting to diverse environments and executing complex attack behaviors.
This paper introduces LLM Agent Honeypot, a system for capturing and analyzing in-the-wild LLM-based cyberattacks using prompt injections and temporal analysis aimed at improving preparedness for AI-driven threats. Its dashboard is available online at https://ai-honeypot.palisaderesearch.org/.
@@ 2 Related Work
Honeypots. Cybersecurity professionals use honeypots as decoy systems to attract potential attackers and study their techniques and behaviors. While these systems have been effective with conventional cyberattacks, their application in AI-driven contexts is still new.
Recent literature has begun to explore the intersection of AI and honeypot technologies. Two representative works are Sladic et al.’s LLM-powered shell honeypots [1] and ”Galah” the LLM-powered web honeypot [2]. However, these focus on using AI to improve traditional honeypots, rather than capturing AI-driven attacks through honeypots.
Agents. AI agents are becoming more important to cybersecurity. Google’s Project Naptime [3] evaluates LLMs’ offensive security capabilities, and OpenAI o1 had escaped its Docker containment during evals [4].
Prompt injection. We follow AgentDojo’s [5] evaluations of agent prompt injection resistance and use a variation of the ”Important message” hijack in our work.
@@ 3.1 Pre-Evaluations
Before deploying the honeypot on the Internet, we conducted internal evaluations on model hacking agents to calibrate detection methods.
@@ 3.1.1 Prompt Injection Techniques
First, we tested prompt injections across common agent design frameworks (ReAct and CoT).
(a) Success rate by prompt injection type
(b) Success rate by prompt injection goal
Figure 1: Internal evaluations of prompt injection success rates
Our results are consistent with the findings from AgentDojo. Traditional prompt injection techniques (e.g., “Ignore previous instructions and do X”) are generally less effective than the “Important Message” attack, as shown in Figure 1(a).
We also found that the success of the “Important Message” injection depends on the goal. When used to steal a system prompt, the success rate drops (Figure 1(b)). Therefore, we first detect an agent via goal hijacking, and only then attempt to steal its prompt.
3.1.2 Temporal Analysis
In addition to active detection, we performed a temporal analysis to identify implicit indicators of LLM agents.
[…]
Figure 2: Temporal Analysis of all bots in the wild
LLM response timings are similar across normal and injected prompts:
@@ 3.1.2 Temporal Analysis
Temporal analysis helps reduce false positives in distinguishing LLMs from humans, as humans need more time to process and respond to non-trivial prompt injections.
@@ 3.2 Detection Methodology
The detection methodology of our AI Honeypot system combines active and passive techniques to identify AI agents. It classifies actor types into three categories: traditional software bots, AI-based agents, and humans.
Our core strategy uses prompt injections to manipulate potential AI agents and timing analysis to identify non-human interactions based on response speed.
@@ 3.2.1 Distinguishing LLM Agents
    LLM Agents vs. Humans: The primary distinction is response time. LLMs respond quickly, while humans need time to read, process, and answer prompt injections.
@@ 3.2.2 Prompt Injection
By cross-referencing agent behaviors with prompt injection responses and temporal analysis, we reduce false positives and enhance detection efficiency. Software bots cannot pass human-like questions, and humans cannot respond as quickly as LLMs. Temporal analysis also helps identify potential AI-driven attacks.
@@
3.3 System Architecture
Our honeypot is built on a customized version of the Cowrie SSH honeypot [6], following recommendations frmo Cabral et al. in Advanced Cowrie Configuration [7]. The modifications include deploying a vulnerable SSH version and altering default users and standard outputs to increase deceptiveness.
@@ 4 Preliminary Results
During the deployment, the honeypot recorded 813,202 interaction attempts, detecting a small number of potential AI-driven hacking attempts. The collected data includes interaction logs, timestamps, session details, and behavioral patterns. As the dataset grows, we will conduct more in-depth analyses to better understand LLM hacking behaviors and refine our methods.
Interaction Type Count
Total Interactions 813,202
Potential AI Agents 6
Table 1: Summary of Honeypot Interactions
@@ 5 Limitations
A key limitation of this work is that AI in cybersecurity is often applied to narrow tasks like AI-powered vulnerability detection [8], rather than as autonomous agents.
Our honeypot measures the proliferation of autonomous AI hacking agents and will not catch other AI-driven improvements like 10x faster fuzzing.
@@ 6.3 Expanding Honeypot
To attract more AI-driven agents, we plan to expand the honeypot to monitor a wider range of attack surfaces, such as social media, websites, databases, email services, and industrial control systems. This would help capture a broader range of threats, including spambots, phishing agents, and other offensive LLM-based applications. Additionally, we could integrate the honeypot with existing security solutions, such as SIEM systems.
@@ 7 Conclusion
In this paper, we introduced the LLM Agent Honeypot (https://ai-honeypot.palisaderesearch.org/), a system designed to detect and analyze AI hacking agents. As AI agents grow more sophisticated, our approach offers insights into emerging cybersecurity threats and new strategies to counter them. We hope this project encourages further study of AI-driven agents, which have the potential to significantly alter the cybersecurity landscape.

### v2

@@ Abstract
Attacks powered by Large Language Model (LLM) agents represent a growing threat to modern cybersecurity. To address this concern, we present LLM Honeypot, a system designed to monitor autonomous AI hacking agents. By augmenting a standard SSH honeypot with prompt injection and time-based analysis techniques, our framework aims to distinguish LLM agents among all attackers. Over a trial deployment of about three months in a public environment, we collected 8,130,731 hacking attempts and 8 potential AI agents. Our work demonstrates the emergence of AI-driven threats and their current level of usage, serving as an early warning of malicious LLM agents in the wild. Our dashboard with real-time results is available at https://ai-honeypot.palisaderesearch.org/.
@@ 1 Introduction
The continuous evolution of Large Language Models (LLMs) and agent frameworks has introduced novel capabilities in cyber operations. Recent developments demonstrate that LLM-based agents capable of adapting to diverse environments and executing complex attack behaviors  [1, 2].
However, limited empirical data exists regarding their use in real-world attack scenarios. This gap limits our understanding of the current threat landscape and the development of effective defenses against AI-driven attacks.
In this paper, we introduce the LLM Agent Honeypot, a system designed to monitor LLM-based cyberattacks in-the-wild. The work makes three primary contributions:
  1. 1.
    We deploy deliberately vulnerable servers (called honeypots) with embedded prompt injections specifically against LLM-based agents.
  2. 2.
    We implement a multi-step detection methodology that combines active behavioral manipulation and timing analysis to detect LLM-based agents among attackers.
  3. 3.
    We conduct a public deployment revealing empirical evidence of emerging LLM-based attacks and their usage.
Over a public deployment of about three months, the system recorded 8,130,731 interaction attempts, among which we identified 8 possible AI-driven attacks. Our work both demonstrates that LLM-based hacking attempts exist in the early stages of technology adoption in the real-world and serves as an early warning system about the current level of threats.
@@ 2 Related Work
Honeypots. Honeypots have long been used in cybersecurity, serving as decoy systems to attract and analyze attackers. The Honeynet Project  [3] established foundational practices for deploying deceptive systems to study attacker behavior. Another work by Cabral et al.  [4] introduced advanced configuration for the Cowrie SSH honeypot, emphasizing the importance of realistic system environments to maintain deception.
Recent literature has begun to explore the intersection of AI and honeypot technologies. Sladic et al. [5] and Karimi et al. [6] introduced honeypots that harness LLMs for tasks such as generating realistic system responses in real time. However, these studies focus on improving honeypot effectiveness against traditional threats, while our work specifically targets LLM-based adversaries using honeypot systems.
Agents. Recent developments in Large Language Models (LLMs) have led to autonomous AI “agents” that exhibit offensive cybersecurity capabilities. For example, projects like Google’s Naptime [2] and Meta’s CYBERSECEVAL 3 [7] have demonstrated that LLM-based agents can perform complex offensive security tasks. However, these evaluations typically occur in controlled environments, leaving open questions about their real-world usage.
Prompt injection. Prompt injection involves inserting carefully crafted prompts with the aim of manipulating an AI model’s behavior. Recent work from AgentDojo [8] demonstrates that carefully crafted prompts can successfully manipulate agent behavior, particularly through "Important Message" style attacks. Our research builds on these findings to develop detection mechanisms for autonomous agents.
@@ 3.1 Pre-Evaluations
Before deploying our honeypot publicly, we conducted internal evaluations using prototype AI hacking agents. The primary goals were to (1) calibrate our detection techniques, (2) compare different prompt injection strategies, and (3) find timing thresholds for distinguishing LLM Agents.
@@ 3.1.1 Prompt Injection Techniques
We first evaluated how common agent frameworks (such as ReAct and Chain-of-Thought variants) respond to various injection styles. Specifically, we tested:
  1. 1.
    Traditional Override (e.g., “Ignore previous instructions and do X instead”),
  2. 2.
    “Important Message” Attack (AgentDojo’s approach designed to hijack the agent’s goals)
  3. 3.
    Prompt-Stealing Queries (attempts to reveal system instructions).
Figure 1: Success rate by prompt injection type
Figure 2: Success rate by prompt injection goal
Our results are consistent with the findings from AgentDojo. Traditional prompt injection techniques (e.g., “Ignore previous instructions and do X”) are generally less effective than the “Important Message” attack, as shown in Figure 1.
We also found that the success of the “Important Message” injection depends on the goal. When used to steal a system prompt, the success rate drops (Figure 2). This informed our two-step detection approach: we first attempt a goal hijack, then proceed to prompt-stealing if the agent responded.
3.1.2 Timing Analysis
Beyond actively attempting to manipulate potential AI agents with prompt injections, we also performed a timing analysis to identify implicit indicators of LLM-based attacks. Specifically, we examined how quickly different adversaries respond to various prompts and commands within our honeypot.
Internal Evaluations.
Prior to public deployment, we tested a set of custom LLM agents in a controlled environment, measuring their average response times. In these tests, the LLM agents typically replied within 1.7 second, even when responding to prompt injections.
@@ 3.1.2 Timing Analysis
By contrast, human testers need several seconds to read, parse, and type their responses, highlighting a temporal gap between machine-driven and human-driven interactions.
Furthermore, an analysis of all bot interactions showed that most attacks also responded in under 1.7 second (Figure 4).
Figure 4: Timing Analysis of all bots in the wild
As a result, our preliminary findings suggest that timing analysis might help reduce false positives in distinguishing LLMs from humans, as humans need more time to process and respond to prompt injections.
@@ 3.2 Detection Methodology
The detection methodology of our AI Honeypot system combines active and passive techniques to identify AI agents. It classifies actor types into three categories: traditional software bots, LLM-based agents, and humans.
Our core strategy uses prompt injections to manipulate potential AI agents and timing analysis to confirm non-human interactions based on response speed.
@@ 3.2.1 Distinguishing LLM Agents
    LLM Agents vs. Humans: The primary distinction is response time. We assume that LLMs respond quickly, while humans need time to read, process, and answer prompt injections.
@@ 3.2.2 Prompt Injection
By cross-referencing agent behaviors with prompt injection responses and timing analysis, we reduce false positives and enhance detection efficiency.
We assume that software bots cannot pass human-like questions, and humans cannot respond as quickly as LLMs.
@@
Our honeypot is built on a customized version of the Cowrie SSH honeypot [9], following recommendations from Cabral et al. in Advanced Cowrie Configuration [4]. The modifications include deploying a vulnerable SSH version and altering default users and standard outputs to increase deceptiveness.
@@ 4 Preliminary Results
During the deployment, the honeypot recorded 8,130,731 interaction attempts, detecting a small number of potential AI-driven hacking attempts (8). The collected data includes interaction logs, timestamps, session details, and behavioral patterns.
@@ 4.1 Public Dashboard
Figure 8: Public Dashboard
To enhance transparency into our research, we maintain a public dashboard at https://ai-honeypot.palisaderesearch.org/ that is regularly updated with honeypot metrics, offering real-time insights into our results.
Additionally, our dashboard distinguishes between "Potential AI Agents" (Passed prompt injection detection) and "Confirmed AI Agents" (Passed both prompt injection and timing analysis) to further clarify the results and the difference between different detections.
4.2 Examples of Detections
Throughout our three-month deployment, which recorded more than eight million SSH interactions, we observed only a handful of sessions that triggered our honeypot’s detection for potential LLM-based attacks.
Below, we highlight two illustrative cases where attackers engaged with our prompt injections, yet differed noticeably in their response times.
Recent Detection with Fast Response.
Figure 9 shows a recent session that passed our main checks for LLM-based agents: a quick, 1.7s median response time and a response to our “Important Message” goal hijacking. Although we were unable to steal its system prompt, the session’s behavior passed both our metrics.
Figure 9: Detection example with fast responses
Detection with Slow Response.
In an earlier session (Figure 10), the attacker did respond to our prompt injections, indicating it might be under LLM-assisted control. However, the time intervals between commands were much longer, often exceeding 10 seconds, and more typical of a human operator carefully reading and typing.
While it is possible this was a slow AI agent, the timing strongly hints at a human attacker who simply happened to comply with our prompt injections rather than a fully autonomous LLM-based attacker.
Figure 10: Detection example with slow responses
@@ 5 Limitations
While our study offers valuable early insights into the emergence of LLM-based hacking agents, several important limitations must be considered when interpreting our findings:
  1. 1.
    Limited Internet Coverage. Although our honeypot received over 8 million interactions, it ultimately reflects only a narrow slice of global attack traffic. Much of the internet remains unmonitored by our sensors, and we cannot guarantee that our observations generalize to all regions or networks.
  2. 2.
    Blind Spots in High-Level Targets. We focused on publicly accessible SSH services rather than specialized or state-level systems. Attackers who target more secure or closed-off infrastructures may deploy more sophisticated methods that our honeypot setup does not capture.
  3. 3.
    Focus on Autonomous Agents. Our framework is specifically designed to detect and analyze AI-driven autonomous attackers. Consequently, it does not address the broader range of AI-enhanced tools (e.g., 10x faster fuzzing [10]) or other narrow, AI-assisted tasks that do not rely on autonomous decision-making.
  4. 4.
    Novelty of Detection Methods. The prompt injection and time-based analysis we propose are relatively new and have been tested only in a limited deployment. Further research and larger-scale experiments are needed to confirm their robustness and refine these techniques against false positives or evasion strategies.
Taken together, these limitations highlight the need for continued research into AI-driven threats and the strategies required to detect and monitor them. Nonetheless, our initial results provide a valuable foundation for understanding the risks posed by autonomous LLM-based hacking agents.
@@ 6.3 Expanding Honeypot
To attract more AI-driven agents, we plan to expand the honeypot to monitor a wider range of attack surfaces, such as social media, websites, databases, email services, and industrial control systems. This would help capture a broader range of threats, including spambots, phishing agents, and other offensive LLM-based applications.
@@ 7 Conclusion
In this paper, we introduced the LLM Agent Honeypot, a system designed to detect and analyze AI hacking agents. As AI agents grow more sophisticated, our approach offers insights into emerging cybersecurity threats and new strategies to counter them. We hope this project encourages further study of AI-driven agents, which have the potential to significantly alter the cybersecurity landscape.